# Topological observables and $\theta$ dependence in high temperature QCD from lattice simulations

### v1

@@ 1 Introduction
It all makes the study of topological properties of QCD a very important and interesting problem. One of the particular quantities of interest is the topological susceptibility $\chi$ in QCD at high temperature. At large enough temperatures one could expect that its behaviour is described by the so-called dilute instanton gas approximation (DIGA) Gross:1980br , while deeply in the confinement phase it should approach a non-zero value, $\chi^{1/4}=75.5(5)$ MeV GrillidiCortona:2015jxo , thus solving the $U_{A}(1)$ axial problem. The behaviour of $\chi$ around the phase transition as well as the approach to the DIGA behaviour:
$\chi(T)\propto{T^{\beta_{1}-4}}\prod_{i=1}^{N_{l}}\frac{m_{i}}{T}$ (1)
with $\beta_{1}=11N_{c}/3-2N_{l}/3$, and $N_{c},N_{l}$ number of colors and light flavors, quark masses $m_{i}$, are not yet fully understood. Due to the intrinsically non-perturbative nature of QCD, one of the most reliable approaches to this problem is lattice calculation. However, lattice studies of different groups within different approaches Chen:2022fid ; Athenodorou:2022aay ; Petreczky:2016vrs ; Borsanyi:2016ksw still do not quantitatively agree with each other, requiring further study of this problem.
Beyond the topological susceptibility, the $\theta$-dependence in QCD is relevant both for theoretical and phenomenological reasons Vicari:2008jw . The $\theta$-term and the related chiral phase of the quark masses are collected in the effective vacuum angle $\theta=\theta+\text{Arg\, det}\,M$, where $M$ is the quark mass matrix, which violates CP-symmetry. Exceptions are not only $\theta=0$ (or if one of the quark masses would be zero), but also $\theta=\pi$: in these cases the CP symmetry is exact. However, at $\theta=\pi$ there may be degenerate vacuum states with a possible spontaneous breaking of CP symmetry Witten:1979vv ; Witten:1980sp ; Smilga:1998dh ; Vonk:2019kwv . In addition, the $\theta$ dependence may be a probe of non-perturbative dynamics at finite temperature, including the fate of the CP breaking transition and its possible endpoint. However so far this aspect has been investigated mostly in model studies Gaiotto:2017yup ; Bigazzi:2015bna .
@@
On the lattice, in full QCD only the first cumulant $\chi$ and the second cumulant $b_{2}$ have been computed Bonati:2015vqz ; Burger:2018fvb , and we will present below our results for physical pion mass and Wilson fermions. These are important probes of the dynamics, but their knowledge does not allow a full reconstruction of the Free Energy. Analytic continuation is an option (we remind the reader that due to the sign problem a direct calculation at non-zero $\theta$ is not possible), but only a limited set of results is available, mostly in Yang-Mills theories Bonanno:2024ggk ; DElia:2013uaf ; Panagopoulos:2011rb .
@@ 3.2 Chiral susceptiblity
We estimate chiral condensate and chiral susceptibility using 24 stochastic sources.
@@ 3.4 Renormalization for Wilson twisted mass fermions
Fermionic method for the topological susceptilibility (18) uses the lattice results for the disconnected chiral susceptibility. The chiral observables (6)–(7) get additional renormalization. The chiral condensate $\bar{\psi}\psi$ has both additive and multiplicative renormalization factors. Leading order additive divergence is proportional to $\mu/a^{2}$. In the disconnected chiral susceptibility additive divergences are subtracted out and thus are absent. Since combinations $m_{l}\bar{\psi}\psi$ and $m_{l}^{2}\chi$ (after subtracting the additive divergences) are RG-invariant Aoki:2006we , the multiplicative renormalization factor for the chiral condensate is equal to the inverse mass renormalization factor $Z_{m}^{-1}$, and for the chiral susceptibility (both connected and disconnected) it equals $Z_{m}^{-2}$. $Z_{m}$ is related to the renormalization constant of the pseudoscalar density operator $Z_{p}$: $Z_{m}=Z_{p}^{-1}$ Frezzotti:2004wz . The values of $Z_{p}$ for ensembles studied in the current work were calculated during zero temperature simulations by the ETM Collaboration in ExtendedTwistedMass:2022jpw , and we present them in Tab. 2. In our analysis, in agreement with Eq. (19), we use pseudoscalar renormalization factor $Z_{p}\left[\leavevmode\nobreak\ \overline{MS}(2\leavevmode\nobreak\ \mathrm{GeV})\leavevmode\nobreak\ \right]$ calculated in $\overline{MS}$-scheme at the scale $\mu=2$ GeV.
@@ 5 Results on topological susceptibility from fermionic method
In Fig. 5 we plot the same results on a double log scale. The plot shows three different temperature regions: around the critical temperatures, up to about 250 MeV; crossover region between 250 and 300 MeV; and a high temperature regime. The straight lines shown in the plot are fits to the DIGA inspired form $\chi^{1/4}_{\mathrm{top}}\propto T^{-2}$, and at this stage are meant to guide the eye. Note that the $b=2$ in Eq. (1) corresponds to $N_{l}=3$, while for $N_{l}=2$, $b=1.91$. The dashed dotted lines represent fits with the logistic curve in the crossover region.
@@ 5.1 Continuum limit
Since we perform the simulations in the fixed scale approach, i.e. only for fixed temperature values for each ensemble, we need to interpolate the data for intermediate temperature values in order to perform the continuum extrapolations at fixed temperature. We used two different ways to interpolate: a simple linear interpolation between two consecutive points and a monotonic cubic interpolation, with the difference between two methods included in the systematic uncertainty. Results of the interpolation (with both systematic and statistical uncertainties) are presented by bands of the corresponding color in Fig. 4.
@@ 5.4 DIGA regime
For the behaviour on individual ensembles, fits with an open exponent extrapolates to a continuum value consistent either with $b=(1.91,2)$, corresponding to $N_{l}=2,3$ respectively, Fig. 7. Alternatively, in the Fig. 5 we have shown fits with a DIGA exponent constrained to the $N_{l}=3$ value. In this region the continuum extrapolation gives reasonable results, as we have discussed. In Fig. 8 we show these fits, together with the extrapolated band.
@@ 6.4 Towards the Grand Canonical Partition Function
$\begin{split}Z(\theta,T)=\sum_{n}Z_{Q}(T)e^{iQ\theta},\\ f(\theta,T)=-\frac{1}{V}\frac{Z(\theta,T)}{Z(0,T)}.\end{split}$ (23)
@@
At zero temperature Chiral Perturbation Theory predicts the well know $\cos(\theta/2)$ dependence GrillidiCortona:2015jxo for the energy. In the same paper, the temperature corrections have been estimated as well, see Fig. 15. The main effect when temperature increases is the modification of the cusp at $\theta=\pi$: the derivative around $\theta=\pi$ gets smaller. As the cusp is associated with a discontinuity of the topological charge, the transition weakens with temperature. It is well known, since the early work by Gasser and Leutwyler Gasser:1987ah , that chiral perturbation theory breaks down at temperature of around $100$ MeV — this is natural from the perspective of an effective theory, where the symmetry pattern plays an important role. Anyway this interval suffices to appreciate the deformation of the energy, confirmed by the consistence of LO and NLO corrections in Fig. 15. (If, just as an exercise, we extend the results of ChPT at higher temperature we observe instabilities at $T=300$ MeV.)
@@ 6.5 Results
To monitorize the behaviour of the zeros of the GCPF, we follow $Z(\theta=\pi)/Z(\theta=0)$, see Fig. 19. What we observe is that $Z(\theta=\pi)/Z(\theta=0)$ is ’very small’ (perhaps compatible with zero) up to $T\simeq 300$ MeV. It would be very tempting to interpret the ’small’ $Z(\pi)$ as a smoking gun for a nearby complex zero, evolving towards real zero in the infinite volume limit, as predicted by Ref. Akemann:2001ir . In this spirit, the observed $T=300$ MeV threshold may be inteprested as the threshold beyond which the Dashen phenomenon disappears – but, admittedly, all this is a wishful thinking as some guidance from analytic calculations would be desirable. At this stage, we can say that our results indicate a significant change of the properties around $\theta=\pi$ at temperatures of about $300$ MeV.
@@ 7 Summary and conclusions
We presented the results of our lattice study of the QCD topological properties at high temperatures $T>180$ MeV, carried out using $N_{f}=2+1+1$ Wilson twisted mass fermions at the maximal twist with the physical value of the pion mass. We generated ensembles at three different lattice spacings and studied the continuum limit.
We presented the topological susceptibility determined from the fermionic method, using the disconnected chiral susceptibility, and the gluonic method, using the field-theoretical definition of the topological charge density, which was measured on the smeared configurations with the Wilson flow. We confirm that the latter definition has larger cutoff effects, and, consequently, larger final uncertainty. Within this largish uncertainty, we have found agreement between the topological susceptibility determined from both gluonic and fermionic methods on our ensembles, see Fig. 13. Note that the agreement between gluonic- and fermionic-based definitions of topological charge at high temperatures was also observed with staggered discretization in Ref. Petreczky:2016vrs with a pion mass of $160$ MeV. Similar results were obtained using spectral projectors methods in Ref. Athenodorou:2022aay , again with staggered fermions. Earlier studies with Wilson fermions have found agreement between gluonic and fermionic methods on a limited set of temperatures and larger quark masses Burger:2018fvb . The present study has explored a large set of temperatures ranging from $180$ MeV to $600$ MeV with a physical pion mass, using Wilson Fermions, with the inclusion of a dynamical charm. Earlier results with the same action were limited to larger pion masses Burger:2018fvb .
@@
Our results suggest three different regimes: low-temperature with $T<250$ MeV, fast and narrow crossover around $T\sim 300$ MeV and consistent with Dilute Instanton Gas Approximation behaviour above $T\gtrsim 300$ MeV. DIGA behaviour has been consistently observed in the temperature dependence of the topological susceptibility and of the kurtosis $b_{2}$. Moreover, the exploratory calculation of the Free Energy afforded for the first time the direct observation of a $\cos(\theta)$ behaviour in the same range of temperature $T>300$ MeV. The fast and narrow crossover — seen in the topological susceptibility, in $b_{2}$, in the Free Energy $f(\theta,T)$ and the partition function $Z(\pi)/Z(0)$ — happens in the same region where our previous study Kotov:2021rah suggest the limit of the scaling window for the transition. Other studies have associated this threshold to deconfinement, or to new pattern of symmetries. From a theoretical viewpoint, a near-zero value of $Z(\pi)$ may be suggestive of a nearby zero of the $Z$ in the complex plane, hence the growth of $Z(\pi)/Z(0)$ at $T\simeq 300$ MeV. Effective models Akemann:2001ir computed explicitly the zeros in the complex plane at zero temperature, and discuss their relation with spontaneous CP violation and the occurrence of a first order phase transition at $\theta=\pi$, and it may be tempting to associate with the threshold at 300 MeV.

### v2

@@ 1 Introduction
It all makes the study of topological properties of QCD a very important and interesting problem. One of the particular quantities of interest is the topological susceptibility $\chi$ in QCD at high temperature. At large enough temperatures one could expect that its behaviour is described by the so-called dilute instanton gas approximation (DIGA) Gross:1980br , while deeply in the confinement phase it should approach a non-zero value, $\chi^{1/4}=75.5(5)$ MeV GrillidiCortona:2015jxo , thus solving the $U_{A}(1)$ axial problem. Instanton gas at high temperatures is expected to lead to nontrivial chiral effects, pronounced in the spectrum of the Dirac operator Kovacs:2023vzi . The behaviour of $\chi$ around the phase transition as well as the approach to the DIGA behaviour:
$\chi(T)\propto{T^{4-\beta_{1}}}\prod_{i=1}^{N_{l}}\frac{m_{i}}{T}\propto T^{-4b}$ (1)
with $\beta_{1}=11N_{c}/3-2N_{l}/3$, and $N_{c},N_{l}$ number of colors and light flavors, quark masses $m_{i}$, are not yet fully understood. Note that within DIGA temperature dependence of the susceptibility $\chi$ is described by the exponent $b$, which is equal to $b=1.92$ for $N_{l}=2$ flavours and $b=2$ for $N_{l}=3$ light flavours. Due to the intrinsically non-perturbative nature of QCD, one of the most reliable approaches to this problem is lattice calculation. However, lattice studies of different groups within different approaches Chen:2022fid ; Athenodorou:2022aay ; Petreczky:2016vrs ; Borsanyi:2016ksw still do not quantitatively agree with each other, requiring further study of this problem.
Beyond the topological susceptibility, the $\theta$-dependence in QCD is relevant both for theoretical and phenomenological reasons Vicari:2008jw . The $\theta$-term and the related chiral phase of the quark masses are collected in the effective vacuum angle $\theta=\theta+\text{Arg\, det}\,M$, where $M$ is the quark mass matrix, which violates CP-symmetry. Exceptions are not only $\theta=0$ (or if one of the quark masses would be zero), but also $\theta=\pi$: in these cases the CP symmetry is exact. However, at $\theta=\pi$ there may be degenerate vacuum states with a possible spontaneous breaking of CP symmetry Witten:1979vv ; Witten:1980sp ; Smilga:1998dh ; Vonk:2019kwv . In addition, the $\theta$ dependence may be a probe of non-perturbative dynamics at finite temperature, including the fate of the CP breaking transition and its possible endpoint. However so far this aspect has been investigated mostly in model studies Gaiotto:2017yup ; Bigazzi:2015bna ; Parnachev:2008fy ; Zhitnitsky:2008ha .
@@
On the lattice, in full QCD only the first cumulant $\chi(T)\equiv c_{2}(T)$ and the second cumulant $b_{2}(T)\equiv\frac{c_{4}(T)}{12c_{2}(T)}$ have been computed Bonati:2015vqz ; Burger:2018fvb , and we will present below our results for physical pion mass and Wilson fermions. These are important probes of the dynamics, but their knowledge does not allow a full reconstruction of the Free Energy. Analytic continuation is an option (we remind the reader that due to the sign problem a direct calculation at non-zero $\theta$ is not possible), but only a limited set of results is available, mostly in Yang-Mills theories Bonanno:2024ggk ; DElia:2013uaf ; Panagopoulos:2011rb .
@@ 3.2 Chiral susceptiblity
We estimate chiral condensate and chiral susceptibility using $N_{\mathrm{stoch}}=24$ stochastic sources following our previous paper Kotov:2021rah . Using already generated data we repeated the analysis with smaller number of stochastic sources $N_{\mathrm{stoch}}<24$ and checked that for $N_{\mathrm{stoch}}=24$ statistical errors are dominated by the gauge noise, making further increase of the $N_{\mathrm{stoch}}$ unnecessary.
@@ 3.4 Renormalization for Wilson twisted mass fermions
Fermionic method for the topological susceptibility (18) uses the lattice results for the disconnected chiral susceptibility. The chiral observables (6)–(7) get additional renormalization. The chiral condensate $\bar{\psi}\psi$ has both additive and multiplicative renormalization factors. Leading order additive divergence is proportional to $\mu/a^{2}$. In the disconnected chiral susceptibility additive divergences are subtracted out and thus are absent. Since combinations $m_{l}\bar{\psi}\psi$ and $m_{l}^{2}\chi$ (after subtracting the additive divergences) are RG-invariant Aoki:2006we , the multiplicative renormalization factor for the chiral condensate is equal to the inverse mass renormalization factor $Z_{m}^{-1}$, and for the chiral susceptibility (both connected and disconnected) it equals $Z_{m}^{-2}$. $Z_{m}$ is related to the renormalization constant of the pseudoscalar density operator $Z_{p}$: $Z_{m}=Z_{p}^{-1}$ Frezzotti:2004wz . The values of $Z_{p}$ for ensembles studied in the current work were calculated during zero temperature simulations by the ETM Collaboration in ExtendedTwistedMass:2022jpw , and we present them in Tab. 2. In our analysis, in agreement with Eq. (19), we use pseudoscalar renormalization factor $Z_{p}\left[\leavevmode\nobreak\ \overline{MS}(2\leavevmode\nobreak\ \mathrm{GeV})\leavevmode\nobreak\ \right]$ calculated in $\overline{MS}$-scheme at the scale $\mu=2$ GeV.
@@ 5 Results on topological susceptibility from fermionic method
In Fig. 5 we plot the same results on a double log scale. The plot shows three different temperature regions: around the critical temperatures, up to about 250 MeV; crossover region between 250 and 300 MeV; and a high temperature regime. The straight lines shown in the plot are fits to the DIGA inspired form $\chi^{1/4}_{\mathrm{top}}\propto T^{-2}$, and at this stage are meant to guide the eye. Note that the $b=2$ in Eq. (1) corresponds to $N_{l}=3$, while for $N_{l}=2$, $b=1.92$. The dashed dotted lines represent fits with the logistic curve in the crossover region.
@@ 5.1 Continuum limit
Since we perform the simulations in the fixed scale approach, i.e. only for fixed temperature values for each ensemble, we need to interpolate the data for intermediate temperature values in order to perform the continuum extrapolations at fixed temperature [^1] As an alternative, one can try to fit data at a fixed lattice spacing with a physically motivated ansatz and perform the continuum extrapolation of the parameters of this ansatz. However, we have found, that it is a nontrivial task to find a suitable ansatz, which works in a wide range of temperatures and gives a reliable continuum extrapolation with small errors, except for DIGA regime $T\in[350:600]$ MeV, see Sec. 5.4.. We used two different ways to interpolate: a simple linear interpolation between two consecutive points and a monotonic cubic interpolation, with the difference between two methods included in the systematic uncertainty. Results of the interpolation (with both systematic and statistical uncertainties) are presented by bands of the corresponding color in Fig. 4.
@@ 5.4 DIGA regime
For the behaviour on individual ensembles, fits with an open exponent extrapolates to a continuum value consistent either with $b=(1.92,2)$, corresponding to $N_{l}=2,3$ respectively, Fig. 7. Alternatively, in the Fig. 5 we have shown fits with a DIGA exponent constrained to the $N_{l}=3$ value. In this region the continuum extrapolation gives reasonable results, as we have discussed. In Fig. 8 we show these fits, together with the extrapolated band.
@@ 6.4 Towards the Grand Canonical Partition Function
$\begin{split}Z(\theta,T)=\sum_{Q}Z_{Q}(T)e^{iQ\theta},\\ f(\theta,T)=-\frac{1}{V}\log\frac{Z(\theta,T)}{Z(0,T)}.\end{split}$ (23)
@@
At zero temperature Chiral Perturbation Theory predicts the well know $\cos(\theta/2)$ dependence GrillidiCortona:2015jxo for the energy. In the same paper, the temperature corrections have been estimated as well, see Fig. 15. The main effect when temperature increases is the modification of the cusp at $\theta=\pi$: the derivative around $\theta=\pi$ gets smaller. As the cusp is associated with a discontinuity of the topological charge, the transition weakens with temperature. It is well known, since the early work by Gasser and Leutwyler Gasser:1987ah , that chiral perturbation theory breaks down at temperature of around $100$ MeV — this is natural from the perspective of an effective theory, where the symmetry pattern plays an important role. Anyway this interval suffices to appreciate the deformation of the energy, confirmed by the consistence of LO and NLO corrections in Fig. 15. If, just as an exercise, we extend the results of ChPT at higher temperature we observe instabilities at $T=300$ MeV. Similar phenomenon, manifested as negative values of the topological susceptbility, was also observed in GomezNicola:2019myi .
@@ 6.5 Results
To monitor the behaviour of the zeros of the GCPF, we follow $Z(\theta=\pi)/Z(\theta=0)$, see Fig. 19. What we observe is that $Z(\theta=\pi)/Z(\theta=0)$ is ’very small’ (perhaps compatible with zero) up to $T\simeq 300$ MeV. It would be very tempting to interpret the ’small’ $Z(\pi)$ as a smoking gun for a nearby complex zero, evolving towards real zero in the infinite volume limit, as predicted by Ref. Akemann:2001ir . In this spirit, the observed $T=300$ MeV threshold may be interpreted as the threshold beyond which the Dashen phenomenon disappears – but, admittedly, all this is a wishful thinking as some guidance from analytic calculations would be desirable. At this stage, we can say that our results indicate a significant change of the properties around $\theta=\pi$ at temperatures of about $300$ MeV.
@@ 7 Summary and conclusions
We presented the results of our lattice study of the QCD topological properties at high temperatures $T>180$ MeV, carried out using $N_{f}=2+1+1$ Wilson twisted mass fermions at the maximal twist with physical values of the quark masses. We generated ensembles at three different lattice spacings and studied the continuum limit.
We presented the topological susceptibility determined from the fermionic method, using the disconnected chiral susceptibility, and the gluonic method, using the field-theoretical definition of the topological charge density, which was measured on the smeared configurations with the Wilson flow. We confirm that the latter definition has larger cutoff effects, and, consequently, larger final uncertainty. Within this largish uncertainty, we have found agreement between the topological susceptibility determined from both gluonic and fermionic methods on our ensembles, see Fig. 13. Note that the agreement between gluonic- and fermionic-based definitions of topological charge at high temperatures was also observed with staggered discretization in Ref. Petreczky:2016vrs with a pion mass of $160$ MeV. Similar results were obtained using spectral projectors methods in Ref. Athenodorou:2022aay , again with staggered fermions. Earlier studies with Wilson fermions have found agreement between gluonic and fermionic methods on a limited set of temperatures and larger quark masses Burger:2018fvb . The present study has explored a large set of temperatures ranging from $180$ MeV to $600$ MeV at the physical point, using Wilson Fermions, with the inclusion of dynamical strange and charm quarks. Earlier results with the same action were limited to larger pion masses Burger:2018fvb .
@@
Our results suggest three different regimes: low-temperature with $T<250$ MeV, fast and narrow crossover around $T\sim 300$ MeV and consistent with Dilute Instanton Gas Approximation behaviour above $T\gtrsim 300$ MeV. DIGA behaviour has been consistently observed in the temperature dependence of the topological susceptibility and of the kurtosis $b_{2}$. Moreover, the exploratory calculation of the Free Energy afforded for the first time the direct observation of a $\cos(\theta)$ behaviour in the same range of temperature $T>300$ MeV. The fast and narrow crossover — seen in the topological susceptibility, in $b_{2}$, in the Free Energy $f(\theta,T)$ and the partition function $Z(\pi)/Z(0)$ — happens in the same region where our previous study Kotov:2021rah suggests the limit of the scaling window for the transition. Other studies have associated this threshold to deconfinement, or to new pattern of symmetries.
From the viewpoint of the phase diagram in the ($\theta,T$) plane, the observed near-zero value of $Z(\pi)$ at a temperature $T$ may be suggestive of a nearby complex zero of $Z$. Effective models Akemann:2001ir computed explicitly the zeros of $Z$ in the complex plane at zero temperature, and discuss their relation with spontaneous CP violation and the occurrence of a first order phase transition at $\theta=\pi$. Then, it may be tempting to associate the growth of $Z(\pi)/Z(0)$ at $T\simeq 300$ MeV with the endpoint of this first order transition line, although of course further work, and in particular a finite size scaling analysis, would be needed to put this statement on firm grounds.